# From Optimization to Prediction: Transformer-Based Path-Flow Estimation to the Traffic Assignment Problem

### v1

@@ 5.2 Model evaluation indicator
To evaluate the performance of the model, four measures are considered, one based on the UE condition and three classical measures. Based on the first condition of UE principle (Equation 1), we calculate the difference of average delay (AD) of the network between the predicted flow and the solution obtained from the optimizer as follows:
@@
In other words, a smaller difference, ideally close to zero, indicates a better-performing model. Other metrics to evaluate model performance include: mean absolute error (MAE), mean absolute percentage error (MAPE):
@@ 5.3.1 Data description
For the experiment, incomplete OD demand values are achieved by randomly removing a ratio of OD pairs. Let $s$ denotes the number of complete OD pairs $(v1,v2)$, with $v1\neq v2$, $s=N^{2}-N$, $j$ denotes the missing percentage. Two missing ratios are considered in this experiment: $j=30\%$ and $j=40\%$. With this experiment, we can perform ‘what-if’ analyses where the demand changes. To create the training, testing, and validation datasets, we generate 4000 OD demand matrices at each missing ratio for the graph (sjs-j OD pairs). This approach enables the model to learn the inherent patterns and structures of the transportation network, even when some demand information is missing.
@@ 5.3.2 Numerical results
Average path cost (mins) 216.62 387.51
Predicted average delay (mins) 2.64 5.34
Delay percentage (%) 1.22 1.38
AD difference (mins) 2.62 4.61
In addition to the prediction accuracy, the training time is also an important factor in evaluating the efficiency and practicality of machine learning models. The training time is 72 minutes, while the time for solving the UE problem with the optimizer is 210 minutes. The prediction time for a single OD demand matrix is 0.001 seconds, it is 5000 times faster than using the Gurobi optimizer (5 - 7 seconds per OD matrix). These results highlight the computational efficiency of the proposed model in addressing traffic assignment problems.
@@ 5.4 Validation on urban transportation network
Scenario 3: There are multiple classes in the network (car and truck).
@@
Scenario 4: When the network scale increases (large-scale network).
Finally, we test the model on a larger-scale urban network to examine its scalability and robustness in more complex settings. Compared to the synthetic and medium-sized urban networks, this scenario involves a greater number of OD pairs, links, and potential sources of congestion. It allows us to assess the model’s efficiency and effectiveness when applied to realistic, large-scale transportation systems.
@@ 5.4.1 Data description and Experiment Set-up
In scenarios 1, 2, and 3, we utilize Sioux Falls network. The information about this network is sourced from H. et al. (2023). This network comprises 24 nodes and 76 links. Table 6 and Figure 5 illustrates the network’s characteristics.
@@
The original dataset comprises link lengths ranging from 2 to 10 kilometers and free-flow travel times spanning 2 to 10 minutes. The histograms of these two variables exhibit a right-skewed distribution, indicating that the majority of links are shorter in length and have shorter free-flow travel times. Most links have capacities between 5,000 and 10,000, while links with capacities exceeding 20,000 constitute a small percentage of the dataset H. et al. (2023).
@@
Having generated the optimal dataset to serve as the ground truth for performance evaluation, we proceed to train the model to obtain its predictions. The training settings and hyperparameters are kept identical to those used in the previous experiment. The OD demand values and the number of OD matrices are kept constant across all experiments, with $30\%$ of OD entries missing and a total of $4000$ OD matrices generated. Scenarios 1-3 utilize the Sioux Falls network, which consists of 24 nodes and 76 links, while Scenario 4, designed to test scalability, employs the larger Eastern Massachusetts (EMA) network, which includes 74 nodes and 258 links. The experimental setup for each scenario is described as follows.
@@
In Scenario 3, we extend the framework to a multi-class setting with $n=2$ classes (cars and trucks). Truck demand is randomly generated as $50\%$ of the car demand, and the free-flow travel time for trucks is set to $1.5$ times that of cars. This scenario uses the full network with all $L$ links.
Table 7: Eastern-Massachusetts network details
@@
Length (miles) 1.06 32.925
Capacity (vehicle) 825 8352.013
@@
In Scenario 4, we evaluate the scalability of the proposed framework using the larger EMA urban network. Table 7 provides detailed information about the EMA network. As this scenario focuses on assessing the model’s performance with large-scale networks, we conduct a single experiment with the full network, where 50% of the demand is missing.
@@ 5.4.2 Numerical results
Table 8 summarizes the prediction performance of the model under various scenarios of missing links. It is evident that the model can predict path flow distribution even with missing links, although it performs best when no links are missing. The predicted path flows result in a network delay of 0.79% compared to the average path cost, which corresponds to a difference of 0.65 minutes relative to the optimal solution’s delay. The error of predicted path flow is smaller than the error of link flow, which is aggregated from predicted path flow, with MAPE of 2.25% for path flows and 2.61% for link flows.
@@
Predicted average delay (mins) 0.71 1.96 2.15
Delay percentage (%) 0.79 1.52 1.69
AD difference (mins) 0.65 1.90 2.06
@@
Figure 9: Heat map of link flow MAPE in different missing ratios of link
@@
Table 9 presents the model’s performance when two and three links are randomly disabled without retraining. The results demonstrate that the proposed model is capable of accurately predicting the distribution of the path flow despite minor changes in the network supply, without requiring retraining. The predicted path flow exhibits a MAPE of less than 5%, indicating high predictive accuracy with a small error. Furthermore, the average network delay associated with the predicted path flow remains low, at just 0.65%. As the number of removed links increases from two to three, the MAPE of the predicted path flow rises from 3.61% to 4.25%. The AD difference is minimal, showing the model’s strong ability to predict travel times accurately.
@@
Predicted average delay (mins) 0.95 1.05
Delay percentage (%) 0.63 0.65
UE Solution average delay (mins) 0.07 0.06
AD difference 0.88 0.99
@@
Delay percentage (%) 0.36 0.35
UE Solution delay percentage (%) 0.17 0.09
AD difference (mins) 2.35 2.27
Scenario 4:
Table 11 presents the performance of the proposed model when applied to the large-scale EMA network for predicting optimal path flow. The results demonstrate that even with a network of this scale, the model achieves a low error rate, with MAPE of 3.69%. The delay percentage of the network, based on the predicted path flow, is 2.36%, only 0.21% higher than the "ground-truth" solution. This indicates that the model’s predictions are closely aligned with optimal User Equilibrium (UE) conditions in terms of travel time delay. Furthermore, the model’s prediction time for each OD matrix is 0.003 seconds, which is 5000 times faster than the 15-second solving time required by Gurobi.
Table 11: Model performance with EMA network (50% OD missing)
Indicator EMA network (50% OD missing)
MAE 6.01
MAPE (%) 3.69
Delay percentage (%) 2.36
UE Solution delay percentage (%) 2.15
AD difference (mins) 24.72
For this complex problem, utilizing 4,000 samples provides an appropriate balance between avoiding overfitting and maintaining computational efficiency. A smaller sample size may increase the risk of overfitting, while a larger sample size could result in significantly higher computational costs. Given that the current framework is adaptable to both small and large-scale networks, it is not necessary to conduct experiments with a more complex framework.
@@ 6 Discussion and Conclusion
To validate the proposed method, we conducted experiments on both synthetic networks and real-world urban transportation networks, considering various link-missing cases and multi-class scenarios with incomplete OD demand, and on a large-scale network, demonstrating the model’s effectiveness in predicting optimal path flow distributions. The results demonstrate that the model can reliably predict path flows even when both OD demand data and network link data are missing. Moreover, it maintains robust performance under network topology changes without retraining, despite being trained only on baseline missing OD demand. Thus, the framework not only reduces computational costs but also adapts to changes in network demand and structure without retraining, enabling robust ‘what-if’ analyses for transportation planning and management. This study demonstrates the effectiveness of an attention-based sequence learning framework in capturing complex correlations between OD pairs (Figure 1), providing a robust surrogate model capable of accelerating complex optimization tasks in areas such as resource allocation and infrastructure management.
@@
Given the methodological focus of this study, and in view of differences in output representation and conservation treatment across existing models, direct benchmarking against state-of-the-art approaches is not pursued here. Moreover, most existing studies do not provide publicly available code or reproducible implementations, making it challenging to ensure a fair and consistent comparison within the same path-flow framework. A systematic comparative evaluation in a unified path-flow setting remains an important direction for future work. Currently, the model focuses on learning and predicting static traffic flow patterns, with the approach limited to path exploration, cost functions, and a graph-based representation. However, the framework is flexible enough to accommodate additional inputs, such as topological considerations, which could further improve equilibrium predictions. As a potential extension, the framework could be adapted to learn dynamic traffic flow patterns. Future research will explore how this model can be trained to handle time-dependent path flows.

### v2

@@ 5.2 Model evaluation indicator
To evaluate the performance of the model, four measures are considered, one based on the UE condition and three classical measures. Based on the first condition of the UE principle (Equation 1), we calculate the difference of average delay (AD) of the network between the predicted flow and the solution obtained from the optimizer as follows:
@@
In other words, a smaller difference, ideally close to zero, indicates a better-performing model. To address the potential scaling effects of the denominator, we additionally examined the distribution of magnitude of the excess cost relative to mean travel times, per-OD pair in A.
Other metrics to evaluate model performance include: mean absolute error (MAE), mean absolute percentage error (MAPE):
@@ 5.3.1 Data description
For the experiment, incomplete OD demand values are achieved by randomly removing a ratio of OD pairs. Let $s$ denotes the number of complete OD pairs $(v1,v2)$, with $v1\neq v2$, $s=N^{2}-N$, $j$ denotes the missing percentage. Two missing ratios are considered in this experiment: $j=30\%$ and $j=40\%$. With this experiment, we can perform ‘what-if’ analyses where the demand changes. To create the training, testing, and validation datasets, we generate 4000 OD demand matrices at each missing ratio for the graph (sjs-j OD pairs). This approach enables the model to learn the inherent patterns and structures of the transportation network, even when some demand information is missing. For each incomplete-OD scenario, the UE reference solution is computed using the same OD matrix with missing entries (rather than the corresponding complete-demand matrix), and this scenario-consistent UE solution is used as the ground truth for training and evaluation.
@@ 5.3.2 Numerical results
Predicted Delay (%) 1.22 1.38
UE Solution Delay (%) 0.0001 0.0001
In addition to the prediction accuracy, the training time is also an important factor in evaluating the efficiency and practicality of machine learning models. The training time is 72 minutes, while generating the optimization-based reference dataset (i.e., solving the UE problem for all 4000 demand scenarios) required 210 minutes in total. The prediction time for a single OD demand matrix is 0.001 seconds, it is 5000 times faster than using the Gurobi optimizer (5 - 7 seconds per OD matrix). These results highlight the computational efficiency of the proposed model in addressing traffic assignment problems.
@@ 5.4 Validation on urban transportation network
Scenario 3: Generalisation to Topological Variations
In this scenario, we generate a set of synthetic networks by modifying the original SiouxFalls urban network used in the previous experiments. This setup simulates structural changes such as urban expansion or the introduction of new routes resulting from the closure or redevelopment of existing areas. The model is trained on the original SiouxFalls network and tested on these topological variations to assess its generalization capability to unseen network structures.
Scenario 4: There are multiple classes in the network (car and truck).
@@
Scenario 5: When the network scale increases (large-scale network).
Finally, we test the model on two larger-scale urban networks to examine its scalability and robustness in more complex settings. Compared to the synthetic and medium-sized urban networks, this scenario involves a greater number of OD pairs, links, and potential sources of congestion. It allows us to assess the model’s efficiency and effectiveness when applied to realistic, large-scale transportation systems.
@@ 5.4.1 Data description and Experiment Set-up
In scenarios 1, 2, 3 and 4 we utilize Sioux Falls network. The information about this network is sourced from Bar-Gera et al. (2023). This network comprises 24 nodes and 76 links. Table 6 and Figure 5 illustrates the network’s characteristics.
@@
The original dataset comprises link lengths ranging from 2 to 10 kilometers and free-flow travel times spanning 2 to 10 minutes. The histograms of these two variables exhibit a right-skewed distribution, indicating that the majority of links are shorter in length and have shorter free-flow travel times. Most links have capacities between 5,000 and 10,000, while links with capacities exceeding 20,000 constitute a small percentage of the dataset Bar-Gera et al. (2023).
@@
Having generated the optimal dataset to serve as the ground truth for performance evaluation, we proceed to train the model to obtain its predictions. The training settings and hyperparameters are kept identical to those used in the previous experiment. The OD demand values and the number of OD matrices are kept constant across all experiments, with $30\%$ of OD entries missing and a total of $4000$ OD matrices generated. Scenarios 1-3 utilize the Sioux Falls network, which consists of 24 nodes and 76 links, while Scenario 4, designed to test scalability, employs the larger Eastern Massachusetts (EMA) network, which includes 74 nodes and 258 links and the Anaheim Network, which includes 38 zones covering 416 nodes and 914 links. The experimental setup for each scenario is described as follows.
@@
In Scenario 3, we generate a set of synthetic networks by modifying the original SiouxFalls urban network used in the previous experiments. Specifically, 20 unique topological variants are created for three configurations of link-change fractions: 25%, 50%, and 75%. Each variant is formed by randomly adding and removing links, with the feature values of new links computed as the mean of their neighbouring links to ensure realistic network characteristics. Figure 10 illustrates three representative randomly generated network variations corresponding to each case. The model is trained on the original SiouxFalls network and tested on these variants to assess its robustness under structural variations.
In Scenario 4, we extend the framework to a multi-class setting with $n=2$ classes (cars and trucks). Truck demand is randomly generated as $50\%$ of the car demand, and the free-flow travel time for trucks is set to $1.5$ times that of cars. This scenario uses the full network with all $L$ links.
In Scenario 5, we evaluate the scalability of the proposed framework using the larger EMA and Anaheim urban network. Table 7 provides detailed information about the EMA and Anaheim networks. For the Anaheim network, number of zones is not same as number of nodes, so for each path connecting the zones, the link features are computed as the average of the links comprising the connecting path. Additionally, to generate OD variations, we create multiple trip demand scenarios using a scaling factor on the base trip demand for Anaheim, resulting in 4000 OD variations. As this scenario focuses on assessing the model’s performance with large-scale networks, we conduct a single experiment with the full network, where 50% of the demand is missing.
Table 7: Network details for EMA and Anaheim
(a) Eastern Massachusetts (EMA) network
@@
Length (miles) 1.06 32.93
Capacity (vehicles) 825 8352.01
@@
(b) Anaheim network
Detail Min value Max value
Length (feet) 264 9451
Capacity (vehicles) 1800 12600
Free flow travel time (mins) 0.05 3.57
Demand 1 2106
@@ 5.4.2 Numerical results
Table 8 summarizes the prediction performance of the model under various scenarios of missing links. It is evident that the model can predict path flow distribution even with missing links, although it performs best when no links are missing. The predicted path flows result in a network delay of 0.79% compared to the average path cost.The error of predicted path flow is smaller than the error of link flow, which is aggregated from predicted path flow, with MAPE of 2.25% for path flows and 2.61% for link flows.
@@
Prediction Delay (%) 0.79 1.52 1.69
UE Solution Delay (%) 0.02 0.03 0.05
@@
For the full Sioux Falls network with 0% missing links, the path-level MAE of 5.02 corresponds to a link-level MAE of 256.27. This larger absolute error at the link level is expected because link flows are obtained by aggregating multiple path flows through the path-link incidence matrix. Therefore, even small errors at the path level may accumulate on links that are shared by many OD pairs, especially high-demand or frequently used links.
The error distribution in Figure 8 should therefore be interpreted in light of the heterogeneous usage of links and paths across the network. Most predicted errors remain small, while larger errors are concentrated on links or paths carrying higher demand or serving as common segments for several OD pairs. From a practical perspective, the predicted path flows remain useful because they preserve route-level information and allow link flows to be recovered through aggregation. This is particularly important for planning and scenario analysis, where route-choice patterns and relative flow redistribution are often more informative than isolated link-level errors.
@@
To highlight the proposed model’s accuracy and computational efficiency, we compare its performance against both traditional mathematical methods and a learning-based baseline. Classical traffic assignment algorithms, including MSA, Frank-Wolfe-type algorithms, and reduced-gradient approaches, have been widely studied and benchmarked in the TAP literature (Perederieieva et al., 2015; Babazadeh et al., 2020). In this study, Table 9 presents a comparison with the Method of Successive Averages (MSA) (Robbins, 1951; Bar-Gera, 2002), the Gap Heuristic proposed by Lu et al. (2009), and a Multi-Layer Perceptron (MLP) baseline. The MLP is implemented as a fully connected neural network with four layers of sizes [256, 128, 64, 32]. All methods are evaluated on the full Sioux Falls network with 30% missing OD demand.
The MSA solves the TAP by iteratively shifting a fraction of traffic from non-shortest to shortest paths, where at each iteration $i$ the swapped fraction is fixed as $\frac{1}{i+1}$. In the gap-based variant, this fraction is adaptively determined from the relative gap between path flows at each iteration, allowing larger shifts when the system is far from equilibrium and smaller adjustments as the gap decreases. The high delays observed for MSA stem from its heuristic, iterative nature (Lu et al., 2009; Ameli et al., 2020a). In early iterations, all demand is concentrated on a few shortest paths, and even after many iterations, Gap-based MSA remains far from equilibrium, resulting in inflated average delays.
Structurally, in the MLP model, the OD demand data and network data are flattened into a single vector as input. Thus, each OD pair is treated independently. This formulation lacks the ability to model interdependencies between ODs. As a result, it produces reasonable estimates at the individual OD level but fails to enforce network-wide consistency. In other words, the flattened representation of MLP does not explicitly encode the structured relationships among OD pairs, paths, and network topology. Instead, all dependencies must be learned implicitly through dense layers, without the benefit of attention-based interaction modeling. This structural disadvantage causes MLP to perform worse than MSA despite being trained on exact Gurobi UE solutions. In contrast, MSA inherently preserves these dependencies through its iterative link flow updates. In comparison, Gurobi-based optimization provides near-optimal solutions by explicitly solving the TAP as a convex optimization problem. However, as noted previously, optimization-based models require recalculating the entire network for even minor ’what-if‘ scenarios.
The Transformer model, trained on Gurobi-generated UE path flows, achieves high accuracy while offering substantially faster inference than both optimization-based frameworks and traditional heuristic approaches, as indicated by the computation time. It is important to note that the computation time reported for MSA corresponds to a solution that is still far from equilibrium, whereas, within the same time frame, the optimization-based framework attains near-optimal results. A trained Transformer model, on the other hand, predicts path flows roughly 1000 times faster than either approach for this network. The MLP predicts approximately 30 times faster than the Transformer due to its significantly simpler architecture.
Table 9: Average delay (in percentage) and computation time of different methods on the original network (0% missing links), sorted by performance
Method Average Delay (%) Computation Time (s)
Gurobi Optimizer 0.03% 3
Transformer Prediction Model 0.79% 0.003
Gap-Based MSA 4.56% 4.2
MLP 15.16% 0.0001
@@
Figure 9: Heat map of link flow MAPE in different link missing ratios
@@
Table 10 presents the model’s performance when two and three links are randomly disabled without retraining. The results demonstrate that the proposed model is capable of accurately predicting the distribution of the path flow despite minor changes in the network supply, without requiring retraining. The predicted path flow exhibits an MAPE of less than 5%, indicating high predictive accuracy with a small error. Furthermore, the average network delay associated with the predicted path flow remains low, at just 0.65%. As the number of removed links increases from two to three, the MAPE of the predicted path flow rises from 3.61% to 4.25%. The AD difference is minimal, showing the model’s strong ability to predict travel times accurately.
@@
Prediction Delay (%) 0.63 0.65
UE Solution Delay (%) 0.04 0.03
@@
Table 11 presents the model’s performance under varying levels of topological variation, where the model is trained solely on the original SiouxFalls urban network. The results show that, although MAE and MAPE increase as the degree of structural change grows, the predicted delay percentage remains relatively stable across all scenarios. This indicates that while path-flow predictions become less precise when the network structure deviates, the resulting deviations relative to overall network travel times are minimal, demonstrating robust predictive performance under network variations.
(a) 25% variation
(b) 50% variation
(c) 75% variation
Figure 10: Illustrations of sampled generalized topological variations of the SiouxFalls network, showing increasing levels of structural change. Link colors indicate network modifications: red links correspond to removed links, while green links correspond to added links.
Table 11: Model performance under different link-change fractions in topological variations
Link-change fraction MAE MAPE (%) Predicted Delay (%) UE Solution Delay (%)
25% 17.11 3.53 0.60% 0.03%
50% 15.38 3.13 0.50% 0.03%
75% 31.06 5.74 0.64% 0.03%
Scenario 4:
@@
Prediction delay (%) 0.36 0.35
UE Solution delay (%) 0.17 0.09
Scenario 5:
Table 13 presents the performance of the proposed model when applied to the large-scale EMA and Anaheim network for predicting optimal path flow. The results demonstrate that even with a network of this scale, the model achieves a low error rate, with MAE of 6.01 and 1.43 respectiveley. The delay percentage of the network, based on the predicted path flow, is 2.36% and 0.36%, only 0.21% and 0.36% higher than the "ground-truth" solution. This indicates that the model’s predictions are closely aligned with optimal UE conditions in terms of travel time delay. Furthermore, the proposed model achieves a prediction time of 0.003 seconds per OD matrix for the EMA network, representing approximately a 5,000-fold reduction compared to the 15-second solution time required by Gurobi. For the Anaheim network, the model attains a prediction time of 0.0017 seconds per OD matrix, which is more than 20,000 times faster than the 36-second solving time obtained using Gurobi.
Table 13: Model performance comparison between EMA and Anaheim networks (50% OD missing)
Indicator EMA Network Anaheim Network
MAE 6.01 1.43
MAPE (%) 3.69 16.92
Prediction delay percentage (%) 2.36 0.36
UE Solution delay percentage (%) 2.15 0.0001
For this complex problem, utilizing 4,000 samples provides an appropriate balance between avoiding overfitting and maintaining computational efficiency. A smaller sample size may increase the risk of overfitting, while a larger sample size could result in significantly higher computational costs.
@@ 6 Discussion and Conclusion
To validate the proposed method, we conducted experiments on both synthetic networks and real-world urban transportation networks, considering various link-missing cases and multi-class scenarios with incomplete OD demand, and on a large-scale network, demonstrating the model’s effectiveness in predicting optimal path flow distributions. The results demonstrate that the model can reliably predict path flows even when both OD demand data and network link data are missing. Moreover, it maintains robust performance under network topology changes without retraining, despite being trained only on original urban network with baseline missing OD demand. Thus, the framework not only reduces computational costs but also adapts to changes in network demand and structure without retraining, enabling robust ‘what-if’ analyses for transportation planning and management. This study demonstrates the effectiveness of an attention-based sequence learning framework in capturing complex correlations between OD pairs (Figure 1), providing a robust surrogate model capable of accelerating complex optimization tasks in areas such as resource allocation and infrastructure management.
@@
Given the methodological focus of this study, and the differences in output representation and flow conservation treatment across existing models, a direct benchmark against state-of-the-art approaches was not pursued. At equilibrium, link flows are unique, but the corresponding path flows are not. Extending link-based models to predict path flows by modifying their final layer would not yield a uniquely defined or comparable reference, as multiple feasible path-flow sets can correspond to the same UE link-flow distribution. Such an extension would also require solving a linear program over all possible paths, leading to high computational complexity as network size grows. Moreover, most existing studies do not provide publicly available code or reproducible implementations, making it challenging to ensure a fair and consistent comparison within the same path-flow framework. A systematic comparative evaluation in a unified path-flow setting remains an important direction for future work.
We note that as the number of OD pairs increases, both memory and computational requirements grow substantially due to the model operating over all OD-path combinations. In such large-scale settings, a practical solution is zone-based aggregation, where individual OD pairs are grouped into higher-level zones. This reduces the effective number of OD interactions and provides a coarser but computationally tractable representation of the network. There is therefore an inherent trade-off between spatial resolution and computational scalability. Recent work has begun to address these scalability challenges directly. Liu et al. (2026, 2025b) decompose large-scale network problems into parallel subproblems solved efficiently via GPU-accelerated solvers, demonstrating that significant computational gains are achievable without sacrificing solution quality. Integrating such approaches with the proposed framework represents a promising direction for future work.
Currently, the model learns and predicts static traffic flow patterns over a predefined path set and is not designed to handle scenarios where the predefined path set becomes infeasible due to structural network changes, such as link closures or capacity disruptions. While training on augmented disruption scenarios is a conceptually straightforward extension, it introduces a significant data generation burden, as each disruption scenario requires regenerating feasible path sets and re-solving the traffic assignment problem to obtain ground truth flows. This makes large-scale augmentation computationally expensive. Consequently, a path-agnostic encoding based on intrinsic path features such as travel time, number of links or path length among others represents a more practical and scalable architectural direction, as it may improve robustness to path set changes without requiring additional ground truth data. As another potential extension, the framework could be adapted to predict dynamic traffic flow patterns by incorporating time-dependent path flows.